Chapter 07 Group work - STU.xlsx Problem 4a Page 10 On January 1, 2020, Beerbos ACCOUNTS RECEIVABLE balance totalled $40,0
Answer #1

Based on the information provided in the question, answers are as under:

1. Beerbo's beginning accounts receivable (A.R.) balance was $ 40,000

2. Beerbo's ending A.R. balance was $ 20,000

ending A.R. balance= (opening A.R. balance + Sales on account - Collections)

=40,000 + 360,000 - 380,000 = $ 20,000

3. Beerbo's average A.R. balance was $ 30,000

Average A.R. balance = ((opening A.R balance + closing A.R. balance) /2 )

=((40,000+ 20,000)/2) = $ 30,000

4. Beerbo's 2020 sales on account was $ 360,000

5. Beerbo's approximate 2020 A.R. turnover was 12

A.R. turnover = Net credit sales / Average A.R.

= $ 360,000 / 30,000 = 12

6. In 2020 it took Beerbo approximately how many days to collect its outstanding A.R. ?

days to collect outstanding A.R.= 365 / A.R. turnover ratio

= 365 / 12

= 30 days ( 30.42 days rounded to nearest day)

In 2020 it took Beerbo approximately 30 days to collect its outstanding A.R.